Science Report – May 04th

Here is the Science Report for 4 May 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No further emergency evacuation locations were scouted.

2. Sunspot and Solar Flare Monitoring: One solar flare was spotted today on the sun’s southwestern side.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working and were not taken on EVA today.

4. GPS Route Measurements: Both EVAs today took the GPS with them to track their routes for their PHEnOM colleague.

5. Waterless dish cleaning: Bacteria growth and monitoring came to an end today. The study showed promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: Final measurements were taken for all crew members during EVA today. Analysis of the data will occur after the mission is completed.

Crew 193 XO and Science Officer

Copyright © The Mars Society. All rights reserved. | Main Site